Diffusion Morning report 01/21/03

ted berg tberg at snf.stanford.edu
Wed Jan 21 06:29:39 PST 2004


Good Morning all,
    Looking good this morning. SiGe has a zero drift on N2/ar we will 
adjust.  TEOS2 is moving along. Colby is still taming THERMCOs
waiting for Coral names for THERMCOs.  EVerything else OK.

Diffusion 	Date: 	1/21/2004 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-Dirty Ox 	down 	TEOS Conversion 	 
Tylan 8-SiGe 	down 	leak n2/ar 	.6/30
Tylan 9-Poly 	Up 	  	23.7/30
Tylan 10-Nitride 	Up 	  	22.8/30
Tylan 11-BPSG-LTO 	Up 	  	14.8/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	17.5/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	 running 	 
Tylan 16 	N/A 	  	 
Burn Box 	Up 	  	 
ZYGO 	Up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well 	Currently 280- B(c4) 
on line


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20040121/e9ab927c/attachment.html>


More information about the diffusion mailing list